# Dynamic asymptotic dimension growth for group actions and groupoids

### v1

@@ 2.1. Equivalent definitions
which is an irreducible cover of $X$. As $m_{R}(\mathcal{U})\leq m_{R}(\mathcal{V})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$, $m_{R}(\mathcal{U})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$ ( here the equality can be obtained ).
Step 2. In this step, we build $\widetilde{\textnormal{ad}_{X}}(R)+1$ boxes. Since $m_{R}(\mathcal{U})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$ and the equality can be obtained, there exists $x\in X$ such that $B(x,R)$ intersects $\widetilde{\textnormal{ad}_{X}}(R)+1$ elements $U_{0},\cdots,U_{\widetilde{\textnormal{ad}_{X}}(R)}$ of $\mathcal{V}$. Thus $d(U_{i},U_{j})\leq 2R$, for all $i,j\in\{0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)\}$ with $i\neq j$, and we put $U_{0},\cdots,U_{\widetilde{\textnormal{ad}_{X}}(R)}$ into $\widetilde{\textnormal{ad}_{X}}(R)+1$ different boxes.
Step 3. In the final step, we put the remaining sets of the cover $\mathcal{U}$ into these $\widetilde{\textnormal{ad}_{X}}(R)+1$ boxes. Set $\mathcal{U}\big{\backslash}\bigcup\limits_{i=0}^{\widetilde{\textnormal{ad}_{X}}(R)}U_{i}=\{U_{k}\}_{k\in I}$, and we denote by $\Box_{l}$ the $l$-th box. We next claim that for any $x\in U_{k}$, there exists $l\in\{0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)\}$ such that
@@
which implies that $B(y,R)$ intersects $\widetilde{\textnormal{ad}_{X}}(R)+2$ sets $U_{k},B_{0},\cdots,B_{\widetilde{\textnormal{ad}_{X}}(R)}$ of $\mathcal{U}$, which contradicts that $m_{R}(\mathcal{U})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$. Hence, for any $x\in U_{k}$, $x$ must be in some box. Below, we divide $U_{k}$ into several pieces and put them into boxes one by one. Let
@@
Then we put $U^{(l)}_{k}$ into the box $\Box_{l}$ for $l=0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)$ in ascending order. If $U^{(l)}_{k}$ is the empty set, we consider the next box. Observe that $U^{(l)}_{k}$ is a subset of $U_{k}$ and $U_{k}=\bigsqcup\limits_{l=0}^{\widetilde{\textnormal{ad}_{X}}(R)}U^{(l)}_{k}$. Similarly, we can divide all elements of $\mathcal{U}\big{\backslash}\bigcup\limits_{j=0}^{\widetilde{\textnormal{ad}_{X}}(R)}U_{j}$ into $\widetilde{\textnormal{ad}_{X}}(R)+1$ boxes. Therefore, we obtain new boxes and for all $l\in\{0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)\}$,
@@
Let $\mathcal{W}_{l}=\{B_{l}\mid B_{l}\in\Box_{l}\}$, then each $\mathcal{W}_{l}$ is a uniformly bounded, $R$-disjoint family and $X=\bigcup\limits_{l=0}^{\widetilde{\textnormal{ad}_{X}}(R)}\mathcal{W}_{l}$. By definition, it is obvious that $\widehat{\textnormal{ad}_{X}}(R)\leq\widetilde{\textnormal{ad}_{X}}(R)$, which yields $\widehat{\textnormal{ad}_{X}}\preceq\widetilde{\textnormal{ad}_{X}}$.
By the transitivity of the equivalence relation, we obtain that $\textnormal{ad}_{X}\approx\widehat{\textnormal{ad}_{X}}$ by Lemma 2.4. ∎
Next, we provide a couple of typical examples of groups and spaces with asymptotic dimension growth. We mainly consider four classes of growth types: exponential growth, subexponential growth, polynomial growth, and slow growth. Say that a monotone increasing function $f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$ has subexponential growth if for any $b>1$,
@@
Observe that if $f$ grows polynomially, it obviously grows subexponentially, since polynomial functions grow slower than exponential functions. Say that $X$ has polynomial (resp. subexponential, exponential) asymptotic dimension growth, if there exists a polynomial (resp. subexponential, exponential) function $f$ such that $\textnormal{ad}_{X}\preceq f$. Recall from [Dra00] that $X$ has slow asymptotic dimension growth, if there exists a linear function $f$ such that $\textnormal{ad}_{X}\preceq f$.
Like the volume growth, a finitely generated group has at most exponential asymptotic dimension growth.
@@
For a finitely generated group $\Gamma$ with the word metric, there exists $\alpha>0$ such that $ad_{\Gamma}(R)\preceq e^{\alpha R}$.
In [Oza12], Ozawa showed that a metric space with subexponential asymptotic dimension growth has property A. Therefore, the following two examples have property A.
@@
The following example tells us that given a strictly increasing function $f$, we can construct a topological space of asymptotic dimension function $f$. This provides an approach to build spaces with slow asymptotic dimension growth when $f$ grows slowly.
@@ 2.2. The case of coarse spaces
In this subsection, we specialize in the case of coarse spaces. In the case of a coarse space for which the coarse structure is induced from the metric, we observe that $R$-disjointness and uniformly boundedness can be converted into $E$-separatedness and $E$-boundedness in coarse geometry. Further, we show that statements about finite asymptotic dimensions on metric spaces and coarse spaces are equivalent. Moreover, we will define the asymptotic dimension function for such coarse spaces, and prove that the function is the same as the function defined in Proposition 2.5. In the rest of this paper, we will study separable coarse spaces.
@@
Below, we consider the special case of the coarse structure coming from a metric. Let $(X,d)$ be a discrete metric space. For $R>0$, we define a controlled tube
@@
Then the controlled sets are defined as subsets $E$ of $X\times X$ contained in some controlled tube $E_{R}$ for some $R$, which form a coarse structure on $X$.
From the perspective of coarse geometry, we can correspond $R$-disjointness and uniformly boundedness to $E$-separatedness and $E$-boundedness in the following definition.
@@
For the finite asymptotic dimension, the statements on the coarse space and the metric space are equivalent in the above special case.
@@
Remark 2.14.
We can complete the proof by the method analogous to the next lemma.
In the case of coarse spaces, we give an alternative definition of the asymptotic dimension function equal to the function defined in Proposition 2.5. This definition plays an important role in the proof of Theorem 4.12 in Section 4.
Lemma 2.15.
$\mathcal{U}=\mathcal{U}_{0}\sqcup\cdots\sqcup\mathcal{U}_{m}$
such that each $\mathcal{U}_{i}$ is $E_{R}$-separated. Then $\widecheck{\textnormal{ad}_{X}}=\widehat{\textnormal{ad}_{X}}$, hence ${\textnormal{ad}_{X}}\approx\widecheck{\textnormal{ad}_{X}}$.
Proof.
Given $R>0$. Suppose that $\mathcal{U}_{0},\cdots,\mathcal{U}_{\widehat{\textnormal{ad}_{X}}(R)}$ are uniformly bounded, $R$-disjoint families of subsets of $X$ covering $X$. Since each family $\mathcal{U}_{i}$ is uniformly bounded, the cover $\mathcal{U}=\bigsqcup\limits_{i=0}^{\widehat{\textnormal{ad}_{X}}(R)}\mathcal{U}_{i}$ is uniformly bounded. It follows that there exists a constant $C>0$ such that for each $U\in\mathcal{U}$, $\mathrm{diam}(U)\leq C$, which implies that each set $U\times U$ is contained in a controlled set $E_{C}$, thus the cover $\mathcal{U}$ is $E_{C}$-bounded. As each family $\mathcal{U}_{i}$ is $R$-disjoint, then for every $U\neq U^{\prime}$ in $\mathcal{U}_{i}$, $d(U,U^{\prime})>R$, which implies that $(U\times U^{\prime})\cap E_{R}=\varnothing$, so each $\mathcal{U}_{i}$ is $E_{R}$-separated. Hence, $\widecheck{\textnormal{ad}_{X}}(R)\leq\widehat{\textnormal{ad}_{X}}(R)$, which yields $\widecheck{\textnormal{ad}_{X}}\leq\widehat{\textnormal{ad}_{X}}$.
Conversely, given $R>0$, for the controlled set $E_{R}$, there exist a controlled set $F$ and a cover $\mathcal{U}=\{U_{i}\}_{i\in I}$ of $X$ such that $\mathcal{U}$ is $F$-bounded and such that $\mathcal{U}$ admits a decomposition
such that each $\mathcal{U}_{i}$ is $E_{R}$-separated. Since $F$ is a controlled set, $F\subseteq E_{C}$ for some $C>0$. As $\mathcal{U}$ is $F$-bounded, each set $U_{i}\times U_{i}$ is contained in $E_{C}$, which implies that $diam(U_{i})\leq C$, so $\mathcal{U}$ is a uniformly bounded cover. Since each $\mathcal{U}_{i}$ is $E_{R}$-separated, for every $U\neq U^{\prime}$ in $\mathcal{U}_{i}$, $(U\times U^{\prime})\cap E_{R}=\varnothing$, thus $d(U,U^{\prime})>R$, which implies that each $\mathcal{U}_{i}$ is $R$-disjoint. Therefore, $\widehat{\textnormal{ad}_{X}}(R)\leq\widecheck{\textnormal{ad}_{X}}(R)$, which yields $\widehat{\textnormal{ad}_{X}}\leq\widecheck{\textnormal{ad}_{X}}$. Hence, $\widehat{\textnormal{ad}_{X}}=\widecheck{\textnormal{ad}_{X}}$, which implies that $\textnormal{ad}_{X}\approx\widecheck{\textnormal{ad}_{X}}$ by Proposition 2.5. ∎
@@ 3. Dynamic asymptotic dimension growth for group actions
In this section, we will introduce the dynamic asymptotic dimension function and the dynamic asymptotic dimension growth for group actions on compact spaces. For a nondecreasing function $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$, we define the $h$-BLR action, which is inspired by the work of Bartels, Reich, and Lück [BLR08, Bar16]. Furthermore, we will show that an $h$-BLR action has dynamic asymptotic dimension function at most $h$. Also, when the function $h$ grows slower than $\sqrt{x}$, we will obtain an equivariant condition of the BLR condition, which is similar to the definition of asymptotic dimension growth for groups. Moreover, it bridges the dynamic asymptotic dimension growth for group actions with the asymptotic dimension growth for groups in the sense of coarse geometry in Section 4, which allows us to construct several exciting examples of dynamic asymptotic dimension growth for group actions.
@@ 3.1. The case of group actions on compact spaces
In this subsection, we will introduce the dynamic asymptotic dimension growth for group actions on compact spaces, which is the dynamical version of asymptotic dimension growth. Moreover, we will give some interesting examples.
To introduce our primary definition, we first recall the notion of length functions on discrete groups.
@@
    $\ell(\gamma^{-1})=\ell(\gamma)$ for all $\gamma\in\Gamma$;
@@
    $\{\gamma\in\Gamma\mid\ell(\gamma)\leq R\}$ is finite for all $R\geq 0$.
Using the concept of proper length functions on discrete groups, we can define the dynamic asymptotic dimension growth for group actions, which is a generalization of the finite dynamic asymptotic dimension for group actions.
@@
Let $\Gamma$ be a discrete group, and let $X$ be a compact Hausdorff space. Let $\Gamma\curvearrowright X$ be an action by homeomorphisms, and let $\ell_{\Gamma}:\Gamma\rightarrow[0,\infty)$ be a proper length function. For any $R>0$, let $\textnormal{dad}_{\Gamma\curvearrowright X}(R)$ be the smallest number over $m\in\mathbb{N}$ with the following property: for the finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$, there is an open cover $\{U_{0},\cdots,U_{m}\}$ of $X$ such that for each $i\in\{0,\cdots,m\}$, the set
@@
is finite. We call $\textnormal{dad}_{\Gamma\curvearrowright X}:\mathbb{R}^{+}\rightarrow\mathbb{N}$ the dynamic asymptotic dimension function of $\Gamma\curvearrowright X$, and call the growth type of the function $\textnormal{dad}_{\Gamma\curvearrowright X}$ the dynamic asymptotic dimension growth of $\Gamma\curvearrowright X$.
@@
    If $\textnormal{dasdim}(\Gamma\curvearrowright X)=\infty$, then both sides of Equation 3.1 are infinite.
@@
Combining theorems in Section 4 and examples of groups with subexponential asymptotic dimension growth, we can provide several examples of group actions with subexponential dynamic asymptotic dimension growth.
@@
Let $N$ be a finitely generated nilpotent group, and $\Gamma$ be a finitely generated group with finite asymptotic dimension. Then the wreath product $N\wr\Gamma$ admits a free, minimal action on the Cantor set which has polynomial dynamic asymptotic dimension growth. In particular, $\mathbb{Z}\wr\mathbb{Z}$ admits a free, minimal action on the Cantor set which has polynomial dynamic asymptotic dimension growth.
@@
3.2. The BLR condtions
In this subsection, we will introduce the BLR condition for a group action derived from the work of Bartels, Reich, and Lück on the Farrell-Jones conjecture [BLR08, Bar16]. There are several benefits to considering this notion. Firstly, it is similar to the definition of the amenability of a group action. Secondly, it is equivalent to the asymptotic dimension growth for a group under certain conditions. Finally, it is closely related to the dynamic asymptotic dimension growth for group actions.

### v2

@@ 2.1. Equivalent definitions
which forms an irreducible cover of $X$. As $m_{R}(\mathcal{U})\leq m_{R}(\mathcal{V})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$, we have $m_{R}(\mathcal{U})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$ ( where the equality can be reached here ).
Step 2. In this step, we build $\widetilde{\textnormal{ad}_{X}}(R)+1$ boxes. Since $m_{R}(\mathcal{U})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$ and the equality is achievable, there exists $x\in X$ such that $B(x,R)$ intersects $\widetilde{\textnormal{ad}_{X}}(R)+1$ elements $U_{0},\cdots,U_{\widetilde{\textnormal{ad}_{X}}(R)}$ of $\mathcal{V}$. Consequently, $d(U_{i},U_{j})\leq 2R$, for all $i,j\in\{0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)\}$ with $i\neq j$. We then put $U_{0},\cdots,U_{\widetilde{\textnormal{ad}_{X}}(R)}$ into $\widetilde{\textnormal{ad}_{X}}(R)+1$ distinct boxes.
Step 3. In the final step, we allocate the remaining sets of the cover $\mathcal{U}$ into these $\widetilde{\textnormal{ad}_{X}}(R)+1$ boxes. Set $\mathcal{U}\big{\backslash}\bigcup\limits_{i=0}^{\widetilde{\textnormal{ad}_{X}}(R)}U_{i}=\{U_{k}\}_{k\in I}$, and let $\Box_{l}$ represent the $l$-th box. We next claim that for any $x\in U_{k}$, there exists $l\in\{0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)\}$ such that
@@
which implies that $B(y,R)$ intersects $\widetilde{\textnormal{ad}_{X}}(R)+2$ sets $U_{k},B_{0},\cdots,B_{\widetilde{\textnormal{ad}_{X}}(R)}$ of $\mathcal{U}$, which contradicts that $m_{R}(\mathcal{U})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$. Therefore, for any $x\in U_{k}$, $x$ must belong to some box. Subsequently, we divide $U_{k}$ into several pieces and place them into boxes sequentially. Let
@@
Then we put $U^{(l)}_{k}$ into the box $\Box_{l}$ for $l=0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)$ in ascending order. If $U^{(l)}_{k}$ is the empty set, we proceed to the next box. Observe that $U^{(l)}_{k}$ is a subset of $U_{k}$, and $U_{k}=\bigsqcup\limits_{l=0}^{\widetilde{\textnormal{ad}_{X}}(R)}U^{(l)}_{k}$. Similarly, we can divide all elements of $\mathcal{U}\big{\backslash}\bigcup\limits_{j=0}^{\widetilde{\textnormal{ad}_{X}}(R)}U_{j}$ into $\widetilde{\textnormal{ad}_{X}}(R)+1$ boxes. Consequently, we obtain new boxes and for all $l\in\{0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)\}$,
@@
Let $\mathcal{W}_{l}=\{B_{l}\mid B_{l}\in\Box_{l}\}$, then each $\mathcal{W}_{l}$ is a uniformly bounded, $R$-disjoint family and $X=\bigcup\limits_{l=0}^{\widetilde{\textnormal{ad}_{X}}(R)}\mathcal{W}_{l}$. By definition, it is evident that $\widehat{\textnormal{ad}_{X}}(R)\leq\widetilde{\textnormal{ad}_{X}}(R)$, which yields $\widehat{\textnormal{ad}_{X}}\preceq\widetilde{\textnormal{ad}_{X}}$.
By the transitivity of the equivalence relation, we conclude that $\textnormal{ad}_{X}\approx\widehat{\textnormal{ad}_{X}}$ by Lemma 2.4. ∎
In the second part of this subsection, we present several typical examples of groups and spaces exhibiting asymptotic dimension growth. We primarily focus on four classes of growth types: exponential growth, subexponential growth, polynomial growth, and slow growth. We define a monotone increasing function $f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$ as having subexponential growth if for any $b>1$,
@@
Observe that if $f$ has polynomial growth, it inherently exhibits subexponential growth, since polynomial functions grow at a slower rate than exponential functions. We say that $X$ has polynomial (resp. subexponential, exponential) asymptotic dimension growth, if there exists a polynomial (resp. subexponential, exponential) function $f$ such that $\textnormal{ad}_{X}\preceq f$. Recall from [Dra00] that $X$ has slow asymptotic dimension growth, if there exists a linear function $f$ such that $\textnormal{ad}_{X}\preceq f$.
Similar to volume growth, a finitely generated group has at most exponential asymptotic dimension growth.
@@
For a finitely generated group $\Gamma$ with the word metric, there exists $\alpha>0$ such that $\textnormal{ad}_{\Gamma}(R)\preceq e^{\alpha R}$.
Ozawa has demonstrated that subexponential asymptotic dimension growth implies Yu’s property A [Oza12], thereby identifying two classes of spaces having this property as follows.
@@
The subsequent example demonstrates that given a strictly increasing function $f$, we can create a topological space with asymptotic dimension function $f$. This provides an approach to construct spaces that have slow asymptotic dimension growth when $f$ has slow growth.
@@ 2.2. The case of coarse spaces
Our objective in this subsection is to establish the third equivalent definition of asymptotic dimension growth. To achieve this, we consider a coarse space for which the coarse structure is induced from the metric. We translate into the language of coarse geometry the two conditions discussed in Proposition 2.5 that define asymptotic dimension growth, and we demonstrate their equivalence.
We initially recollect the notion of coarse spaces, referring the reader to the monograph [Roe03] for a comprehensive treatment of coarse spaces.
@@
Below, we examine the special case of the coarse structure derived from a metric. Let $(X,d)$ be a discrete metric space. For $R>0$, we define a controlled tube as follows:
@@
Then, the controlled sets are defined as subsets $E$ of $X\times X$ contained within some controlled tube $E_{R}$ for some $R$, thereby forming a coarse structure on $X$.
The following definition introduces two concepts of $E$-separatedness and $E$-boundedness within the realm of coarse geometry.
@@
By translating the concepts of $R$-disjointness and uniformly boundedness from Proposition 2.5 into the terms of $E_{R}$-separatedness and $F$-boundedness within the framework of coarse geometry, we obtain the following third equivalent definition. This definition will be used in the proof of Theorem 4.16 in Section 4.
Lemma 2.13.
$\mathcal{U}=\mathcal{U}_{0}\sqcup\cdots\sqcup\mathcal{U}_{m}$
and each $\mathcal{U}_{i}$ is $E_{R}$-separated. Then $\widecheck{\textnormal{ad}_{X}}=\widehat{\textnormal{ad}_{X}}$, hence $\widecheck{\textnormal{ad}_{X}}\approx{\textnormal{ad}_{X}}$.
Proof.
Given $R>0$. Suppose that $\mathcal{U}_{0},\cdots,\mathcal{U}_{\widehat{\textnormal{ad}_{X}}(R)}$ are uniformly bounded, $R$-disjoint families of subsets of $X$ covering $X$. Since each family $\mathcal{U}_{i}$ is uniformly bounded, the cover $\mathcal{U}=\bigsqcup\limits_{i=0}^{\widehat{\textnormal{ad}_{X}}(R)}\mathcal{U}_{i}$ is also uniformly bounded. Consequently, there exists a constant $C>0$ such that for each $U\in\mathcal{U}$, $\mathrm{diam}(U)\leq C$, which implies that each set $U\times U$ is contained in a controlled set $E_{C}$, thus the cover $\mathcal{U}$ is $E_{C}$-bounded. Furthermore, as each family $\mathcal{U}_{i}$ is $R$-disjoint, then for every $U\neq U^{\prime}$ in $\mathcal{U}_{i}$, $d(U,U^{\prime})>R$. This ensures that $(U\times U^{\prime})\cap E_{R}=\varnothing$, indicating that each $\mathcal{U}_{i}$ is $E_{R}$-separated. Hence, $\widecheck{\textnormal{ad}_{X}}(R)\leq\widehat{\textnormal{ad}_{X}}(R)$, which yields $\widecheck{\textnormal{ad}_{X}}\leq\widehat{\textnormal{ad}_{X}}$.
Conversely, given $R>0$, for the controlled set $E_{R}$, there exist a controlled set $F$ and a cover $\mathcal{U}=\{U_{i}\}_{i\in I}$ of $X$ such that $\mathcal{U}$ is $F$-bounded and admits a decomposition
such that each $\mathcal{U}_{i}$ is $E_{R}$-separated. Since $F$ is a controlled set, $F\subseteq E_{C}$ for some $C>0$. As $\mathcal{U}$ is $F$-bounded, each set $U_{i}\times U_{i}$ is contained in $E_{C}$, implying that $diam(U_{i})\leq C$, so $\mathcal{U}$ is a uniformly bounded cover. Since each $\mathcal{U}_{i}$ is $E_{R}$-separated, for every $U\neq U^{\prime}$ in $\mathcal{U}_{i}$, $(U\times U^{\prime})\cap E_{R}=\varnothing$, thus $d(U,U^{\prime})>R$, which implies that each $\mathcal{U}_{i}$ is $R$-disjoint. Therefore, $\widehat{\textnormal{ad}_{X}}(R)\leq\widecheck{\textnormal{ad}_{X}}(R)$, which yields $\widehat{\textnormal{ad}_{X}}\leq\widecheck{\textnormal{ad}_{X}}$. Hence, $\widehat{\textnormal{ad}_{X}}=\widecheck{\textnormal{ad}_{X}}$, which implies that $\textnormal{ad}_{X}\approx\widecheck{\textnormal{ad}_{X}}$ by Proposition 2.5. ∎
Based on the preceding results, we can use any one of $\textnormal{ad}_{X}$, $\widetilde{\textnormal{ad}_{X}}$, $\widehat{\textnormal{ad}_{X}}$, or $\widecheck{\textnormal{ad}_{X}}$ to define the asymptotic dimension growth.
By making use of a proof approach similar to the above, we can demonstrate that the statement about finite asymptotic dimension on the metric space is equivalent to the statement on the coarse space as stated in Remark 2.11.
@@ 3. Dynamic asymptotic dimension growth for group actions
In this section, we will introduce the dynamic asymptotic dimension function and the dynamic asymptotic dimension growth for actions of discrete groups on compact spaces. To investigate the dynamic asymptotic dimension growth of group actions, we will define the $h$-BLR action for a non-decreasing function $h$, which draws inspiration from the conditions formulated by Bartels, Lück, and Reich [BLR08, Bar16]. Additionally, we will establish an equivalent definition of the $h$-BLR action under specific assumptions. Furthermore, we demonstrate that an $h$-BLR action has dynamic asymptotic dimension function at most $h$. By applying this result within the realm of coarse geometry, we can obtain Theorems 4.21 and 4.22. This allows us to present several typical examples of subexponential dynamic asymptotic dimension growth in group actions.
@@ 3.1. The case of group actions on compact spaces
In this subsection, we will introduce the concept of dynamic asymptotic dimension growth for actions by discrete groups on compact Hausdorff spaces. This notion extends the concept of finite dynamic asymptotic dimension for such actions. By integrating examples of groups that demonstrate subexponential asymptotic dimension growth from the previous section with Theorem 4.22 discussed in the following section, we can present several intriguing examples of group actions that exhibit subexponential dynamic asymptotic dimension growth.
To introduce our primary definition, we initially revisit the concept of length functions on discrete groups.
@@
    $\ell(\gamma^{-1})=\ell(\gamma)$, for all $\gamma\in\Gamma$;
@@
    $\{\gamma\in\Gamma\mid\ell(\gamma)\leq R\}$ is finite, for all $R\geq 0$.
By applying the concept of proper length functions on discrete groups, we can define the dynamic asymptotic dimension growth for actions of discrete groups on compact spaces, which generalizes the finite dynamic asymptotic dimension for these actions.
@@
Let $\Gamma$ be a discrete group, and let $X$ be a compact Hausdorff space. Let $\Gamma\curvearrowright X$ be an action by homeomorphisms, and let $\ell_{\Gamma}:\Gamma\rightarrow[0,\infty)$ be a proper length function. For any $R>0$, let $\textnormal{dad}_{\Gamma\curvearrowright X}(R)$ be the smallest number over $m\in\mathbb{N}$ with the following property: for the finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$, there exists an open cover $\{U_{0},\cdots,U_{m}\}$ of $X$ such that for each $i\in\{0,\cdots,m\}$, the set
@@
is finite. We call $\textnormal{dad}_{\Gamma\curvearrowright X}:\mathbb{R}^{+}\rightarrow\mathbb{N}$ the dynamic asymptotic dimension function of $\Gamma\curvearrowright X$, and call the growth type of the function $\textnormal{dad}_{\Gamma\curvearrowright X}$ the dynamic asymptotic dimension growth of $\Gamma\curvearrowright X$. If no such function $\textnormal{dad}_{\Gamma\curvearrowright X}$ exists, we say that $\Gamma\curvearrowright X$ has infinite dynamic asymptotic dimension.
@@
    If $\textnormal{dasdim}(\Gamma\curvearrowright X)=\infty$, then both sides of equation (3.1) are infinite.
@@
Remark 3.4.
Dynamic asymptotic dimension has been estimated for actions of discrete groups on compact spaces. For instance, the dynamic asymptotic dimension of an action by an infinite virtually cyclic group on a compact Hausdorff space is always one, provided the action has the marker property [ALSS21]. Furthermore, the dynamic asymptotic dimension of a free action of a countable discrete group on a zero-dimensional second countable compact Hausdorff space is either infinite or equal to the asymptotic dimension of the group [CJMST23]. We suspect this result may be generalized to the growth case, but we do not seriously pursue this. Moreover, for an infinite countable discrete group $\Gamma$ with asymptotic dimension $d$, the dynamic asymptotic dimension of a free, minimal action of $\Gamma$ on the Cantor set is at most $d$ [GWY17]. In the next section, we will extend this result to subexponential dynamic asymptotic dimension growth. Combining this generalization with examples of groups with subexponential asymptotic dimension growth, we can provide several examples of group actions with subexponential dynamic asymptotic dimension growth.
Here is a classic example of group actions with polynomial dynamic asymptotic dimension growth.
@@
Let $N$ be a finitely generated nilpotent group, and let $\Gamma$ be a finitely generated group with finite asymptotic dimension. Then the wreath product $N\wr\Gamma$ admits a free, minimal action on the Cantor set which has polynomial dynamic asymptotic dimension growth. In particular, $\mathbb{Z}\wr\mathbb{Z}$ admits a free, minimal action on the Cantor set which has polynomial dynamic asymptotic dimension growth.
Another example is the slow dynamic asymptotic dimension growth of group actions.
@@
3.2. Bartels-Lück-Reich condtions
The Bartels-Lück-Reich conditions, abbreviated as BLR conditions, originated from the seminal works of Bartels, Lück, and Reich regarding the Farrell-Jones conjecture [BLR08, Bar16]. Subsequently, Guentner, Willett, and Yu in the paper [GWY17] further summarized and developed these conditions to study the finite dynamic asymptotic dimension of group actions, thereby giving birth to the concept of a $d$-BLR action. In this subsection, we aim to extend this notion to the growth case by defining the $h$-BLR action for a nondecreasing function $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$. Furthermore, when the function $h$ grows at a rate slower than $\sqrt{x}$, we establish an alternative equivalent definition for this action, which is intimately connected to the asymptotic dimension growth of a group. Moreover, we demonstrate that an $h$-BLR action has dynamic asymptotic dimension function at most $h$.